OpenStudy (anonymous):

This is to Shravanm1, squareroot of 7-2x-square root x+2=squareroot x+5, using substitution

OpenStudy (anonymous):

\[\sqrt{7-2x}-\sqrt{x+2}=\sqrt{x+5}\]

OpenStudy (anonymous):

Let's see you solve it tough guy...use substitution

OpenStudy (anonymous):

substitution? why?

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

you can't use substitution rani...you only gave me one equation

OpenStudy (anonymous):

well your in "Calculus 3", solve it great one

OpenStudy (anonymous):

i can solve it, just not with substitution. that's not possible with the given information

OpenStudy (anonymous):

then use square the root

OpenStudy (anonymous):

oh really? how so?

OpenStudy (anonymous):

thats better!

OpenStudy (anonymous):

then how come I did it?

OpenStudy (anonymous):

when you square both sides, you get: ((7-2x)-(2)(sqrt7-2x)(sqrtx+2)+(x+2))=x+5

OpenStudy (anonymous):

how'd you do it with substitution rani? show me??

OpenStudy (anonymous):

you did an error, you have to remove the sqrts

OpenStudy (anonymous):

but not bad ;), i believe you

OpenStudy (anonymous):

Nope. Not yet.

OpenStudy (anonymous):

the formula is: (a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

ahhh, good good... you cannot use substitution and your not supposed to remove the sqr roots. lol

OpenStudy (anonymous):

finish it and let's see

OpenStudy (anonymous):

yeah from there you'd move the free expressions to the left, leaving the two square roots, then solve. its a lot of work and i cant really draw it out

OpenStudy (anonymous):

one question, are there square roots on (7-2x) and (x+2)? ;)

OpenStudy (anonymous):

Yes, in the middle term(2ab)

OpenStudy (anonymous):

im pretty sure on the left and right also

OpenStudy (anonymous):

when you square a square root, it becomes normal

OpenStudy (anonymous):

so is it going to be 49-4x?? or 7-2x?